# Contextual Embedding-based Clustering to Identify Topics for Healthcare Service Improvement

### v1

@@ III-D2 Gibbs Sampling Dirichlet Mixture Model (GSDMM)
Alongside LDA, we also implement the GSDMM for compliant topic identification. The integration of GSDMM is particularly significant given the nature of our data: most survey responses were brief, often comprising 1 to 3 sentences [36]. GSDMM is uniquely suited for such short-texts, as it operates under the assumption that a single document predominantly expresses a single topic. This is a notable deviation from LDA’s approach, where a document can encompass multiple topics, making LDA more suitable for longer texts [37].
The implementation of GSDMM involved modeling each document as a mixture of topics, with a single dominant topic assigned to each document. For a given document $d$, the topic $k$ is assigned based on the conditional probability $P(k\mid d)$, which is calculated using Equation 1. In this equation, $n_{k,-d}$ represents the number of documents assigned to topic $k$, excluding the current document $d$; $N_{-d}$ is the total number of documents excluding $d$; $K$ is the total number of topics; is the count of word $n_{k,w}$ w in topic k; $n_{k}$ is the total number of words assigned to topic $k$; and $V$ is the vocabulary size. The hyperparameters $\alpha$ and $\beta$ serve as priors for the document-topic and topic-word distributions, respectively, analogous to their roles in LDA. Through iterative sampling, GSDMM converged to a stable distribution of topics across the documents, effectively capturing the dominant theme in each text fragment.
@@ III-D3 BERTopic
BERTopic [38] is a topic modeling technique that integrates transformer-based embeddings, dimensionality reduction, clustering, and topic representation to extract meaningful insights from textual data. It begins with a transformer embedding model, such as BERT or RoBERTa, to generate high-dimensional contextual embeddings that capture semantic relationships between words and documents. To enhance computational efficiency and clustering performance, UMAP (Uniform Manifold Approximation and Projection) is applied for dimensionality reduction, preserving essential structures in a lower-dimensional space. The reduced embeddings are then clustered using HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise), which dynamically determines the optimal number of topics based on data density, enabling the discovery of latent topics in an unsupervised manner. This method allows BERTopic to automatically generate topics without requiring the user to specify a predefined number. Additionally, BERTopic framework [39] offers the flexibility to manually specify the desired number of topics, allowing for both unsupervised and user-controlled topic generation. Finally, c-TF-IDF (class-based Term Frequency-Inverse Document Frequency) is employed to extract the most representative words for each cluster, improving topic interpretability. Combining these four components, BERTopic provides a scalable and flexible approach to topic modeling, making it particularly effective for analyzing large, unstructured text data. However, BERTopic has some limitations. It assumes each document belongs to a single topic, which may not fully capture multi-topic documents.
@@
To address the unique challenges posed by our small dataset of short-texts, we developed a customized topic modeling pipeline that integrates a customized pre-processing approach, BERT embeddings and k-means clustering. Our topic identification system leverages the Bidirectional Encoder Representations from Transformers (BERT), a transformer-based model pre-trained on a vast corpus of text, to generate rich, contextualized embeddings. These embeddings are particularly effective in capturing the subtle nuances of short-text data, making them ideal for our application. Additionally, BERT embeddings can effectively handle Out-of-Vocabulary (OOV) words, making them particularly useful for dealing with domain-specific terminology. Importantly, we only utilized the embedding layer of BERT while freezing all other layers. This ensured that the model acted purely as a feature extractor, preventing any modifications to its pre-trained weights. As a result, there is no risk of overfitting or catastrophic forgetting, making our approach particularly well-suited for small datasets [40]. We avoid dimensionality reduction methods, as they can remove key linguistic features, especially in low-resource settings where every word holds significant meaning [41]. The high-dimensional embeddings generated by BERT are preserved to retain the full contextual representation.
Our approach began with careful preprocessing of the textual data. Stopwords were removed, and the text was lemmatized to ensure uniformity. Additionally, we extended the stopword list with domain-specific and custom terms to improve data quality. Each preprocessed comment was then encoded into embeddings using the ‘bert-base-uncased’ [^1] https://huggingface.co/google-bert/bert-base-uncased variant of the pre-trained BERT model. This model was chosen for its balance between performance and computational efficiency, as well as its ability to generalize well across diverse datasets. The uncased variant ensured the focus remained on content, minimizing case-sensitive variations, which is particularly advantageous for short-texts.
The generated embeddings served as input for a k-means clustering algorithm, where we experimented with different numbers of clusters ($n$ = 3, 5, 10 etc) to find the optimal grouping for the data. For topic interpretability, we extracted the top $k$ words (e.g., 5,10) from each cluster by analyzing word frequencies in representative comments. To improve topic interpretability, we first selected a set of representative comments for each cluster based on their proximity to the cluster centroid in the BERT embedding space. These comments were then analyzed to extract the most frequently occurring words, ensuring that the topic words accurately reflected the cluster’s semantic meaning. This iterative process allowed us to refine the topic selection and ensure the clusters were meaningful and cohesive.
@@ IV-A Feedback Filtering
Our dataset, encompassing 439 comments, underwent keyword matching, which served as a quantitative measure of the filtering process’s effectiveness. Notably, the matching frequency of keywords within the comments indicates a substantial impact on the accuracy of the filtering mechanism. Initial results without the use of lemmatization presented 89 keyword matches. The introduction of lemmatization, a process facilitating the reduction of words to their base or ‘lemma’ form, slightly increased this match count to 92.
A further improvement is achieved by generating and matching keywords across different POS after lemmatization. This approach propelled the match count to a peak of 102 instances. This increment underscores the pivotal role of nuanced linguistic techniques in enhancing the sensitivity of our keyword-based feedback-filtering approach.
[…]
@@
IV-B Topic Modeling
In their work, R. Egger et al. highlight the importance of selecting the appropriate model based on the nature of the data and emphasize the need for careful interpretation of results [42]. Building on this perspective, our topic modeling efforts focused on finding the ideal balance between coherence and clarity. To ensure a comprehensive evaluation, we assessed our models’ performance using two metrics: the widely used coherence score ($C_{v}$) for topic interpretability and Rank-Biased Overlap (RBO) for evaluating topic diversity.
Coherence score $C_{v}$ quantitatively assesses the coherence of topics. A higher $C_{v}$ score indicates better performance, suggesting that the top words of a topic are not only frequent but also semantically similar. This indicates a clearer and more meaningful topic representation. The $C_{v}$ measure is derived from the Normalized Pointwise Mutual Information (NPMI) [43]. It accounts for the probability of words co-occurring within topics relative to their individual probabilities. The NPMI tends to be particularly attuned to less common words and is well-suited for analyzing small datasets like ours [44]. It can be calculated using the equation 2 [45]. In this equation, $\mathrm{P}\left(w_{i}^{r},w_{i}^{s}\right)$ is the joint probability of the words occurring together, while $\mathrm{P}\left(w_{i}^{r}\right)$ and $\mathrm{P}\left(w_{i}^{s}\right)$ are the individual probabilities of the words. The epsilon $\epsilon$ is a small number added to prevent division by zero. The $C_{v}$ is defined by equation 3. $\gamma$ is a parameter that can be tuned to give more importance to higher PMI values. This tuning allows for a weighted contribution of word pairs based on their PMI.
To evaluate the effectiveness of different topic modeling approaches, we computed the Coherence Score (Cv) for each model across three datasets: Clickbait-title, 20Newsgroup, and Patient Feedback. The datasets were divided into large and small subsets to assess model robustness across different data scales. Given that the Patient Feedback dataset consists of short-text comments, we tested our models on public short-text datasets for validation. For the analysis, we primarily focus on short-text, especially for small datasets, to evaluate how well our models handle limited data. Coherence scores are measured using the top 10 topic words for large-text datasets and the top 5 topic words for short-text datasets. The best results for short-text are marked boldface in Tables III and IV.
For large datasets, we generated topics at different levels (5, 10, 20, 30, and 40 topics). As shown in Table III, GSDMM consistently outperformed other models in short-text datasets, achieving the highest coherence score of 0.49 for 20 topics in Clickbait-title. This indicates its ability to effectively capture coherent topics in short-texts. Additionally, kBERT also achieved a high coherence score of 0.49 at 30 topics for short-text datasets (Clickbait-title). This motivates further exploration of kBERT for small datasets, where capturing meaningful patterns in limited data is crucial. LDA, while performing well in long-text scenarios, struggled in short-text datasets.
BERTopic, which dynamically determines the number of topics, generated 163 topics in 20Newsgroup (Cv = 0.60) and 178 topics in Clickbait-title (Cv = 0.49). This adaptability allowed BERTopic to excel in large datasets by identifying a diverse range of topics. However, in smaller datasets, it generated a limited number of topics, reducing coherence performance. In the Patient Feedback dataset, BERTopic produced 3 topics with a coherence score of 0.46, indicating moderate effectiveness in short-text scenarios.
For small datasets, we generated 3, 5, 10, 15, and 20 topics. As presented in Table IV, kBERT performed exceptionally well in handling short-text patient feedback, achieving the highest coherence score of 0.53 for 3 topics, demonstrating its strong ability to extract meaningful insights from limited and sparse patient feedback data. Additionally, in the Clickbait-title (500) dataset, kBERT achieved a strong coherence score of 0.45 for 5 topics, reinforcing its strength in short-text small dataset scenarios. This suggests that BERT embeddings are particularly useful for extracting meaningful insights from short and sparse healthcare-related feedback.
[…]
@@ IV-B Topic Modeling
TABLE V: Topics from kBERT with Examples
Topic Name Topic words Example Complain?
Staff Conduct and Cleanliness good, clean, great, everyone, professional Everyone was careful in the emergency department No
Safety Protocols compliance mask, staff, wearing, safe, one The staff’s masks were reassuring in these uncertain times. No
Efficiency of Service doctor, time, appointment, waiting, room Very long wait from the time I entered the room until the doctor arrived Yes
TABLE VI: IRBOavg Score on Top 5 words and 3 Topics
@@
Rank-Biased Overlap (RBO) is a similarity measure for ranked lists that emphasizes top-weightedness and accommodates incomplete or non-conjoint lists [46]. It is particularly useful for evaluating topic diversity in topic modeling by comparing the top-ranked terms across topics. The RBO is defined by the equation 4. $A_{d}$ is the agreement at rank $d$, and $p$ ($10<p<$ 1) adjusts the weight of deeper ranks. Higher $p$ values increase the depth considered, ensuring meaningful topic comparisons. RBO scores range from 0 (no overlap) to 1 (identical lists). Adjust the parameter $p$ to focus on higher or deeper ranks. To provide a more comprehensive diversity measure, we used Inverted RBO $(1-\operatorname{RBO})$, where higher values indicate greater topic diversity by reflecting lower overlap in top-ranked terms. We further compute the Averaged Inverted RBO by averaging pairwise inverted RBO values, as defined by equation 5. Here, A, B, and C represent the ranked lists of top words from three different topics, and their pairwise inverted RBO scores quantify the extent of dissimilarity among them. A higher $\operatorname{IRBO_{avg}}$ value indicates greater diversity in the extracted topics.
In this study, we selected 3 topics for the Patient Feedback dataset and evaluated topic diversity using Average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$). A higher IRBO score indicates greater topic distinctiveness by measuring lower overlap in top-ranked terms. As shown in Table VI, BERTopic and kBERT achieved the highest diversity score (1.00), suggesting completely distinct topic distributions. GSDMM (0.97) also demonstrated strong topic separation, while LDA (0.83) exhibited comparatively lower diversity, indicating some overlap in identified topics.
@@ V Discussion
Our topic modeling analysis identified three primary areas of concern in patient feedback: ‘Staff Conduct and Cleanliness’, ‘Safety Protocols Compliance’, and ‘Efficiency of Service’. Positive feedback highlighted professionalism and hygiene practices, reinforcing confidence in staff attentiveness. Similarly, the emphasis on safety measures, particularly mask compliance, indicated patient reassurance regarding institutional health protocols. However, the most significant complaint centered on ‘Efficiency of Service’, with frequent mentions of long wait times and appointment delays, emphasizing the need for healthcare facilities to enhance operational efficiency. These key areas of focus are summarized in Table V, which presents the dominant topics along with representative terms and examples.
A crucial step in our approach was keyword filtering, which effectively differentiated complaints from non-complaints within the dataset. By applying a targeted keyword-based filtration method, we identified 102 complaint-related sentences out of 439 total comments.
In evaluating the topic modeling approaches, kBERT consistently demonstrated superior performance in handling short-text complaint identification, achieving the highest coherence score ($C_{v}$ = 0.53 for 3 topics) and an IRBO score of 1.00, indicating distinct and well-separated topics. This suggests that BERT-based embeddings effectively capture the nuanced meaning of patient complaints, making it the most reliable model for healthcare-related short-text analysis. GSDMM also performed well, particularly in short-text datasets, as its cluster-based approach assigns a dominant topic to each document, making it effective for structured short-form feedback. BERTopic, while excelling in larger datasets due to its dynamic topic generation, struggled in smaller datasets, as observed in its moderate coherence score ($C_{v}$ = 0.46) for the Patient Feedback dataset. LDA, a traditional topic modeling approach, showed strong performance in long-text settings but struggled with short-text patient complaints, reaffirming its dependence on a larger corpus for meaningful topic extraction.
These findings emphasize the importance of choosing context-aware models like kBERT for short-text analysis in healthcare settings. The combination of keyword filtering and effective topic modeling ensures that patient concerns are identified and addressed accurately, leading to data-driven improvements in service delivery. The insights gained from this study can help healthcare providers prioritize patient concerns, streamline complaint resolution, and enhance overall service efficiency, ultimately improving the patient experience and quality of care.
@@ VI Conclusion
This study explored unsupervised techniques for complaint topic identification in short-text patient feedback, comparing traditional topic modeling methods (LDA, GSDMM) with advanced neural embedding-based approaches (BERTopic, kBERT). Our results demonstrated that kBERT outperforms other models in both coherence and diversity metrics, making it the most effective method for analyzing short and sparse healthcare-related texts. Additionally, a keyword-based filtering approach significantly improved complaint detection. The identified complaint categories—staff conduct and cleanliness, safety protocol compliance, and efficiency of service—provide actionable insights for healthcare service improvement. In the future, we will also explore few-shot and instruction-based prompting to improve complaint classification and topic extraction in healthcare feedback analysis. To achieve this, we will use open-source LLMs (e.g., LLaMA) locally with sufficient computational resources, preventing data leakage and ensuring compliance.

### v3

@@ III-D2 Gibbs Sampling Dirichlet Mixture Model (GSDMM)
Alongside LDA, we also implement the GSDMM for compliant topic identification. The integration of GSDMM is particularly significant given the nature of our data: most survey responses were brief, often comprising 1 to 3 sentences [15]. GSDMM is uniquely suited for such short-texts, as it operates under the assumption that a single document predominantly expresses a single topic. This is a notable deviation from LDA’s approach, where a document can encompass multiple topics, making LDA more suitable for longer texts [11]. The implementation of GSDMM involved modeling each document as a mixture of topics, with a single dominant topic assigned to each document. For a given document $d$, the topic $k$ is assigned based on the conditional probability P(kd)P(k∣d), which is calculated using Equation 1. In this equation, $n_{k,-d}$ represents the number of documents assigned to topic $k$, excluding the current document $d$; $N_{-d}$ is the total number of documents excluding $d$; $K$ is the total number of topics; is the count of word $n_{k,w}$ w in topic k; $n_{k}$ is the total number of words assigned to topic $k$; and $V$ is the vocabulary size. The hyperparameters $\alpha$ and $\beta$ serve as priors for the document-topic and topic-word distributions, respectively, analogous to their roles in LDA. Through iterative sampling, GSDMM converged to a stable distribution of topics across the documents, effectively capturing the dominant theme in each text fragment.
@@ III-D3 BERTopic
BERTopic [10] is a topic modeling technique that combines transformer-based embeddings, dimensionality reduction, clustering, and topic representation to extract insights from text. It uses models like BERT or RoBERTa to generate contextual embeddings, which are reduced in dimensionality using UMAP for efficient clustering. HDBSCAN is then applied to group similar texts and determine the number of topics based on data density. BERTopic also allows users to set a fixed number of topics manually. For interpretability, it employs class-based TF-IDF (c-TF-IDF) to extract representative words per cluster. While scalable and flexible, BERTopic assumes each document belongs to a single topic, which may limit its effectiveness on multi-topic texts.
@@
To address the unique challenges posed by our small dataset of short-texts, we developed a customized topic modeling pipeline that integrates a customized pre-processing approach, BERT embeddings, and k-means clustering. BERT, a transformer-based model pre-trained on a large corpus, generates rich contextual embeddings that capture subtle linguistic nuances and effectively handle Out-of-Vocabulary (OOV) terms, particularly useful for domain-specific language. We only utilized the embedding layer of BERT while freezing all other layers. This ensured that the model acted purely as a feature extractor, preventing any modifications to its pre-trained weights. As a result, there is no risk of overfitting or catastrophic forgetting, making our approach particularly well-suited for small datasets [13]. We avoid dimensionality reduction methods, as they can remove key linguistic features, especially in low-resource settings where every word holds significant meaning [19]. The high-dimensional embeddings generated by BERT are preserved to retain the full contextual representation.
Our approach began with careful preprocessing of the textual data, including stopword removal and lemmatization for consistency. We also extended the stopword list with domain-specific terms to improve data quality. Each comment was then encoded using the ‘bert-base-uncased’ [^1] https://huggingface.co/google-bert/bert-base-uncased model, selected for its balance of performance and generalization. Using the uncased variant helped reduce case-related variability, which is beneficial for short texts. The resulting embeddings were clustered using k-means, with various cluster sizes ($n$ = 3, 5, 10, etc.) tested to find optimal groupings. For interpretability, we identified representative comments closest to each cluster centroid and extracted the most frequent words from them. These words formed the topic labels, ensuring alignment with the semantic meaning of each cluster. This iterative process refined topic selection and enhanced cluster coherence.
@@ IV-A Feedback Filtering
Our dataset, encompassing 439 comments, underwent keyword matching, which served as a quantitative measure of the filtering process’s effectiveness. Notably, the matching frequency of keywords within the comments indicates a substantial impact on the accuracy of the filtering mechanism. Initial results without the use of lemmatization presented 89 keyword matches. The introduction of lemmatization, a process facilitating the reduction of words to their base or ‘lemma’ form, slightly increased this match count to 92. A further improvement is achieved by generating and matching keywords across different POS after lemmatization. This approach propelled the match count to a peak of 102 instances. This increment underscores the pivotal role of nuanced linguistic techniques in enhancing the sensitivity of our keyword-based feedback-filtering approach.
TABLE II: Coherence Score ($C\textsubscript{v}$) on Large Datasets
@@
a generated 178 topics ($C_{v}$=0.48), b generated 163 topics ($C_{v}$=0.60)
TABLE III: Coherence Score ($C\textsubscript{v}$) on Small Datasets
@@
a generated 2 topics ($C_{v}$=0.40), b generated 2 topics ($C_{v}$=0.44)
IV-B Topic Modeling
We aimed to balance coherence and clarity in topic modeling. To evaluate performance, we used two metrics: Coherence score ($C_{v}$) for interpretability and Rank-Biased Overlap (RBO) for topic diversity.
Coherence score ($C_{v}$) quantitatively measures how semantically related the top words in a topic are. A higher $C_{v}$ indicates better topic quality, suggesting that the most representative words frequently co-occur and are contextually meaningful. The $C_{v}$ score is derived from Normalized Pointwise Mutual Information (NPMI) [5], which considers the joint and individual probabilities of word co-occurrence. NPMI is defined in Equation 2, where $\epsilon$ prevents division by zero. The final $C_{v}$ score is calculated as shown in Equation 3, where $\gamma$ adjusts the weight given to high-PMI word pairs.
To assess model effectiveness, we computed the $C_{v}$ coherence score for each method across three datasets: Clickbait-title, 20Newsgroup, and Patient Feedback. These were split into large and small subsets to test robustness at different data scales. As Patient Feedback consists of short-text comments, we included public short-text datasets for validation. Coherence was measured using the top 10 topic words for long texts and the top 5 for short texts. Best-performing short-text results are highlighted in Tables II and III.
For large datasets, we evaluated models across multiple topic counts (5 to 40). As shown in Table II, GSDMM consistently achieved high coherence on short-text data, scoring 0.49 for 20 topics in Clickbait-title. kBERT also performed well, reaching a score of 0.49 at 30 topics, highlighting its strength in small, short-text scenarios. In contrast, LDA struggled with short texts but performed better on long documents. BERTopic, which automatically determines the number of topics, generated 163 topics in 20Newsgroup ($C_{v}$ = 0.60) and 178 in Clickbait-title ($C_{v}$ = 0.49), showing strong performance on large datasets. However, its coherence dropped in smaller datasets, producing only 3 topics ($C_{v}$ = 0.46) in Patient Feedback.
For small datasets, we generated 3, 5, 10, 15, and 20 topics. As shown in Table III, kBERT achieved the highest coherence score of 0.53 for 3 topics on Patient Feedback, demonstrating its effectiveness in short and sparse data. It also performed well on the Clickbait-title (500) dataset, scoring 0.45 for 5 topics, reinforcing the usefulness of BERT embeddings in low-resource short-text scenarios.
Rank-Biased Overlap (RBO) is a similarity measure for comparing ranked lists, placing more emphasis on top-ranked items and handling incomplete or non-conjoint lists [22]. It is particularly useful for evaluating topic diversity by assessing overlap among top topic words. RBO ranges from 0 (no overlap) to 1 (identical lists), with the parameter $p$ controlling the depth of comparison. To better capture topic diversity, we use the inverted RBO $(1-\operatorname{RBO})$, where higher values indicate more distinct topics. We compute the average inverted RBO across all topic pairs, as shown in Equations 4 and 5.
TABLE IV: IRBOavg Score on Top 5 words and 3 Topics
@@ IV-B Topic Modeling
We evaluated topic diversity on the Patient Feedback dataset using Average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$), where higher scores indicate more distinct topics. As shown in Table IV, kBERT and BERTopic achieved perfect diversity scores of 1.00, while GSDMM scored 0.97 and LDA 0.83, indicating some topic overlap in LDA’s results.
@@ V Discussion
Our topic modeling analysis identified three primary areas of concern in patient feedback: ‘Staff Conduct and Cleanliness’, ‘Safety Protocols Compliance’, and ‘Efficiency of Service’. Positive feedback highlighted professionalism and hygiene practices, reinforcing confidence in staff attentiveness. Similarly, the emphasis on safety measures, particularly mask compliance, indicated patient reassurance regarding institutional health protocols. However, the most significant complaint centered on ‘Efficiency of Service’, with frequent mentions of long wait times and appointment delays, emphasizing the need for healthcare facilities to enhance operational efficiency. These key areas of focus are summarized in Table V, which presents the dominant topics along with representative terms and examples. A crucial step in our approach was keyword filtering, which effectively differentiated complaints from non-complaints within the dataset. By applying a targeted keyword-based filtration method, we identified 102 complaint-related sentences out of 439 total comments.
TABLE V: Topics from kBERT with Examples
Topic words Example Complain?
good, clean, great, everyone, professional Everyone was careful in the emergency department No
mask, staff, wearing, safe, one The staff’s masks were reassuring in these uncertain times. No
doctor, time, appointment, waiting, room Long wait from the time I entered the room until the doctor arrived Yes
In evaluating the topic modeling approaches, kBERT consistently outperformed other models in short-text complaint analysis, achieving the highest coherence score ($C_{v}$ = 0.53) and perfect topic diversity (IRBO = 1.00), indicating clear and well-separated topics. This highlights the strength of BERT embeddings in capturing the nuances of patient feedback. GSDMM also performed well, benefiting from its ability to assign a dominant topic per document. In contrast, BERTopic, while effective on large datasets, showed limited performance on smaller ones ($C_{v}$ = 0.46). LDA performed better on long texts but struggled with short, sparse data. These results underscore the value of context-aware models like kBERT in healthcare settings. Combined with keyword filtering, this approach enables accurate identification of patient concerns, supporting data-driven service improvements and enhanced care delivery.
@@ VI Conclusion
This study explored unsupervised techniques for complaint topic identification in short-text patient feedback, comparing traditional topic modeling methods (LDA, GSDMM) with advanced neural embedding-based approaches (BERTopic, kBERT). Our results demonstrated that kBERT outperforms other models in both coherence and diversity metrics, making it the most effective method for analyzing short and sparse healthcare-related texts. Additionally, a keyword-based filtering approach significantly improved complaint detection. The identified complaint categories include staff conduct and cleanliness, safety protocol compliance, and efficiency of service, and provide actionable insights for healthcare service improvement.
Acknowledgment
We gratefully acknowledge the support of the Advancing a Healthier Wisconsin (AHW) for funding this project. The authors used AI tools in this manuscript, including Grammarly for grammar checking and ChatGPT for rephrasing and refining, with careful review to ensure contextual accuracy.